How much work is done when a bucket of mass 1.5 kg with 10 kg of water in it is pulled up from the bottom of a well 8 m deep? Take
g=9:8m=s2
1
Expert's answer
2014-03-29T10:41:13-0400
A = m*g*h = (1.5 +10) * 9.8 * 8 = 901.6 (J)